### v1

@@ 3.4  Comparative statics inside the job (Jevons Paradox)
An increase in skill $s_{i}$ raises $\omega_{i}^{A}$ if and only if the tasks are substitutes in the production function, $\sigma>1$. This is Jevons’ Paradox for knowledge work. Someone skilled at a task will spend more time with it, not less, when demand for that task is elastic. By contrast, when tasks are complement, each task has to be completed to some predefined level, and skilled workers spend less time on satisfying the job requirement. This also highlights that activities, in general, are poor measures of skills. ## Applying the adoption theorems under CES–CET
@@ 4  Discussion and Relation to the Literature
Task-based labor-market models have a long history, and our framework builds on several strands of this literature. To clarify what is new, we organize prior work along five dimensions: (i) job heterogeneity (which task directions matter across jobs); (ii) worker task elasticity $\gamma$ (how flexibly workers reallocate across tasks); (iii) production elasticity $\sigma$ (how tasks combine in output); (iv) skill heterogeneity (how workers differ in task-specific productivity); and (v) technology directionality (whether technical change favors particular task directions).
4.1  Organizing Framework
Our model has four primitives: job requirements $\theta\in\mathbb{R}^{N}_{+}$, worker skills $s\in\mathbb{R}^{N}_{+}$, transformation elasticity $\gamma>0$ (the curvature of the worker’s task-allocation frontier), and substitution elasticity $\sigma>0$ (the curvature of the production aggregator). Technology enters as a direction $t\in S^{N-1}$ and capability $\chi>0$. This nests several important special cases.
Table 1: Task-based models: key parameters
Paper Jobs $\gamma$ $\sigma$ Workers Technology
Welch (1969) NN-dim 0 $F$ NN-dim
Katz and Murphy (1992) NN-dim 0 $F$ NN-dim general
—application 2 types 0 $\sigma$ 2 types factor-augmenting
Autor et al. (2003) 2 tasks $\infty$ 1 1-dim routine-biased
Lazear (2009) 2 skills 0 $\infty$ 2-dim
Acemoglu and Autor (2011) continuum $\infty$ 1 3 types task threshold
Firpo et al. (2011) 5 tasks $\infty$ >1>1 5-dim 5-dim
Cortes (2016) 3 tasks $\infty$ 0 routine-biased
Fortin and Lemieux (2016) NN-dim 0 $\infty$ NN-dim
Acemoglu and Restrepo (2018) continuum $\infty$ $\sigma$ 1 type rotating threshold
Acemoglu and Restrepo (2020) continuum $\infty$ 1 1 type task threshold
Bárány et al. (2020) NN-dim 0 $\infty$ NN-dim
Böhm (2020) $K$ tasks logit $\infty$ routine-biased
—application 3 tasks logit $\infty$ routine-biased
Cavounidis and Lang (2020) NN-dim $\sigma$ $\infty$ NN-dim
Cavounidis et al. (2024) NN-dim $\infty$ $\sigma$ NN-dim within-occ
Autor and Thompson (2025) 1-dim $\infty$ 0 1-dim 1-dim
Freund and Mann (2025) NN-dim $\infty$ 1 NN-dim directional
Althoff and Reichardt (2026) NN-dim $\infty$ $\sigma$ directional
Benzell and Myers (2026) NN-dim 0 $\sigma$ NN-dim 1-dim
Gans and Goldfarb (2026) $n$ tasks $\infty$ 1 1-dim
Bárány and Holzheu (2026) NN-dim 0 quadratic NN-dim
This paper NN-dim $\gamma$ $\sigma$ NN-dim directional
Note: “Jobs” and “Workers” refer to the dimensionality of heterogeneity in job requirements and worker skills, respectively. This is distinct from the number of tasks $N$: a model with many tasks can still be one-dimensional if tasks are ordered by a single attribute (e.g., routine content, expertise level).
Many models take corner cases on at least one of the two curvature margins. It is common to assume either Leontief aggregation ($\sigma=0$) or perfect substitution ($\sigma=\infty$), and either rigid task endowments ($\gamma=0$) or frictionless reallocation across tasks ($\gamma=\infty$). Allowing both $\gamma$ and $\sigma$ to be finite matters because shadow prices and the direction of marginal value depend on both worker flexibility and production complementarity. This is the key step behind our alignment-based adoption condition and the structured intensive margin.
4.2  Multidimensional Heterogeneity
A key feature is that jobs and workers differ along multiple dimensions that cannot be collapsed to a single index. This rules out simple rankings of “high-skill” versus “low-skill” or “routine” versus “nonroutine” as sufficient statistics.
Welch (1969) is among the first to decompose labor into multiple “labor inputs”—what we now call tasks. Workers have fixed endowments of these inputs ($\gamma=0$), and a general production function $F$ aggregates them. This early framework already allows for $N$-dimensional heterogeneity in both jobs and workers.
Firpo et al. (2011) take a step toward higher dimensionality, using five O*NET task measures. Their reduced-form specification writes wages as an inner product of skills and returns—analogous to our $\langle s,\theta\rangle$ structure. Importantly, because task prices vary across occupations, tasks are implicitly not fully traded ($\gamma<\infty$). Their Oaxaca-Blinder decomposition captures changes in task content across occupations, which corresponds to “rotation” in our framework.
Cortes (2016) studies job polarization with three task types and Leontief production ($\sigma=0$). Workers choose tasks freely ($\gamma=\infty$), and routine-biased technical change shifts employment across task categories. Böhm (2020) estimates task prices using a generalized Roy model where workers sort across $K$ tasks via discrete choice (effectively $\gamma=\text{logit}$). Tasks are traded at economy-wide prices ($\sigma=\infty$), and the focus is on recovering the price effects of routine-biased technical change.
Two recent papers treat skills as genuinely multidimensional objects rather than as a scalar index. Bárány et al. (2020) develop and estimate a multi-sector general-equilibrium model of education and sector-occupation choices; jobs are bundles of tasks and workers sort on a multidimensional skill vector that evolves with schooling and experience (Abstract). Bárány and Holzheu (2026) define worker specialization as the distance between a worker’s skill set and those prevalent in the labor market and study its wage-unemployment trade-off in a random-search model with mismatch penalties and skill complementarity (Abstract). Relative to our framework, these papers emphasize schooling, sorting, and search frictions, whereas we focus on within-job task choice and the alignment-based adoption decision.
Benzell and Myers (2026) is among the first to model within-job production with high-dimensional, non-traded tasks. Workers have skill endowments ($\gamma=0$), and tasks combine with CES elasticity $\sigma$. Automation supersedes one task direction. They emphasize that in high dimensions, workers cannot be ranked by a single index. Our model generalizes by allowing $\gamma>0$ (workers can reallocate across tasks) and directional technology.
4.3  Within-Job Task Reallocation
A second feature of our model is within-job task reallocation: workers shift effort across tasks while holding the job fixed. This margin is central to understanding how AI changes work without changing employment.
Lazear (2009) introduces skills with job-specific weights—a direct antecedent of our $(\theta,s)$ formulation. In his baseline model, workers have endowments of two skills ($\gamma=0$), and firms combine them linearly with heterogeneous weights ($\sigma=\infty$). All skills are general, but firms value them differently, generating apparent firm-specific human capital. The model does not include technology, but the skill-weights structure is foundational. Fortin and Lemieux (2016) extend this skill-weights approach to study inequality, combining it with the task framework of Acemoglu and Autor (2011).
Spitz-Oener (2006) provides early measurement evidence that task content changes within occupations over time, with computer use associated with shifts toward nonroutine cognitive tasks. Cavounidis and Lang (2020) microfound dynamic skill formation in a multidimensional setting. In their static model, workers choose task bundles with elasticity $\gamma=\sigma$ (in our notation) and output sums tasks linearly ($\sigma=\infty$), with job-specific task weights corresponding to our $\theta$. Cavounidis et al. (2024) apply this framework empirically, documenting within-occupation shifts in task content (what we call “rotation”) and connecting productivity effects to an envelope condition (our shadow prices).
Autor and Thompson (2025) model expertise as a scalar that determines how task bundles translate into output. Technology compresses the expertise distribution by making low-expertise workers more productive. The task space is implicitly high-dimensional but collapsed to a single index. Our framework preserves the full dimensionality, showing that scalar indices can mask important heterogeneity in alignment.
Freund and Mann (2025) develop a model where workers have multidimensional skill portfolios and jobs bundle multiple tasks, using 38 O*NET task clusters in their application. Tasks are freely tradable ($\gamma=\infty$) and combine with Cobb-Douglas technology ($\sigma=1$). They emphasize alignment between worker skills and job requirements and show that one-dimensional “exposure” measures can miss important heterogeneity. Their wage formula resembles our productivity expression, and their notion of skill-job “correlation” parallels our inner product structure. Our framework generalizes by allowing finite $\gamma$ (non-traded tasks) and finite $\sigma$ (task complementarity).
Althoff and Reichardt (2026) develop a closely related framework where workers allocate time across $N$ tasks ($\gamma=\infty$) and output is CES with elasticity $\rho$. Technical change is task-specific and directional, allowing them to study how AI shifts comparative advantage across occupations. They extend the analysis to occupational choice, showing how task-specific technology shocks induce reallocation across jobs.
Gans and Goldfarb (2026) study an O-ring production function where workers allocate time across $N$ tasks. Automation makes workers redundant in specific tasks, but reallocation occurs within the job. This is close to our setup with $\gamma=\infty$ and $\sigma=1$. The key difference is that their technology is one-dimensional (automation of an adjacent range of tasks), whereas ours is directional (technology produces a bundle of tasks in fixed proportions).
Lin (2025) provides direct evidence for within-job task reallocation in response to AI changes. Using prompt-level data from Midjourney, she shows that most of the output shift after an upgrade comes from users changing their prompts rather than from the capability change itself. This is evidence of rotation in response to changing AI capabilities, exactly the intensive margin our model emphasizes.
4.4  Other Task-Based Models
Several other papers develop task-based frameworks that inform our approach but emphasize different margins.
Katz and Murphy (1992) extend the Welch (1969) framework to study wage inequality: their general model allows $N$ labor inputs with a generic concave production function; their empirical application collapses to two skill types (college/non-college) with CES aggregation ($\sigma\approx 1.4$). They treat technical change as shifts in labor demand without specifying direction.
The skill-biased technical change literature (Katz and Murphy, 1992; Acemoglu, 2002) partitions workers into broad skill groups and models technology as factor-augmenting. This approach captures between-group inequality but cannot address within-occupation heterogeneity in adoption or gains. The routine-biased framework (Autor et al., 2003; Acemoglu and Autor, 2011) adds a task dimension: technology substitutes for routine tasks while complementing non-routine tasks. Workers choose task bundles, but the task space is one-dimensional (routine vs. non-routine), and tasks are traded at economy-wide prices ($\gamma=\infty$).
Acemoglu and Restrepo (2018, 2020) develop a workhorse model where a continuum of tasks is traded at economy-wide prices. Automation shifts the threshold between tasks performed by labor and capital. The production elasticity $\sigma$ governs macroeconomic implications. Our framework nests this as the limiting case $\gamma\to\infty$ (tasks fully tradable) with technology “rotating” the threshold across task indices.
Lise and Postel-Vinay (2020) develop a sorting model with multidimensional skills. Workers and jobs match in a frictional labor market. The focus is on sorting and human capital accumulation rather than technology adoption.
Agrawal et al. (2026) emphasize that AI may relax bottlenecks rather than automate tasks. This is complementary to our approach: in our model, AI that produces tasks where the worker is constrained (high shadow price) relaxes the binding constraint, which is the geometric interpretation of bottleneck relief.
The contribution of our framework is to combine five features that prior models typically treat separately: (i) a high-dimensional task space; (ii) finite worker task elasticity $\gamma$ (non-traded tasks and worker-specific shadow prices); (iii) finite production elasticity $\sigma$ (task complementarity); (iv) multidimensional skill heterogeneity $s$; and (v) directional technology. The interaction of these features yields the adoption cone, the alignment condition, and the intensive margin.
@@ 5  Conclusion
GenAI is an unusually easy-to-adopt general-purpose technology, but its usefulness varies sharply across directions in a high-dimensional task space. This directional heterogeneity makes one-dimensional notions of exposure or skill unreliable guides to either adoption or productivity gains.
We model a job as a neoclassical task allocation problem: workers choose task bundles subject to a convex resource constraint, and the optimal allocation is characterized by a vector of shadow (autarky) prices that summarize which tasks are scarce at the margin in the current worker-job match. We model AI as a directional technology with a capability level. Adoption then reduces to a value comparison at shadow prices.
Three implications follow. First, absolute advantage is necessary but not sufficient: even a superior tool may not be adopted when its strengths point away from what is valuable at the margin. Second, adoption has a sharp entry threshold; near it, small capability gains can produce large expansions in the set of adopted directions. Third, adoption has a structured intensive margin: because entry and all-in thresholds generally differ, optimal use can remain partial over a range of capabilities, generating stable hybrid production.
These mechanisms are match-specific because shadow prices depend jointly on worker skills and job requirements, and they extend naturally to settings with multiple tools, where adoption rotates shadow prices and can raise the bar for subsequent technologies.
@@ A.5  CES-CET closed forms
$\frac{x_{i}}{x_{j}}=\left(\frac{\theta_{i}}{\theta_{j}}\right)^{\sigma/(\gamma+\sigma)}\left(\frac{s_{i}}{s_{j}}\right)^{\gamma/(\gamma+\sigma)}.$
@@
Time shares. From the allocation formula, $\omega_{i}^{A}=x_{i}^{A}/\sum_{j}x_{j}^{A}\propto s_{i}^{\gamma/(\gamma+\sigma)}\theta_{i}^{\sigma/(\gamma+\sigma)}$. Logarithmic differentiation gives $\partial\omega_{i}^{A}/\partial s_{i}=[\gamma/((\gamma+\sigma)s_{i})]\omega_{i}^{A}(1-\omega_{i}^{A})>0$.

### v2

@@ 3.4  Comparative statics inside the job (Jevons Paradox)
An increase in skill $s_{i}$ raises $\omega_{i}^{A}$ if and only if the tasks are substitutes in the production function, $\sigma>1$. This is Jevons’ Paradox for knowledge work. Someone skilled at a task will spend more time with it, not less, when demand for that task is elastic. By contrast, when tasks are complement, each task has to be completed to some predefined level, and skilled workers spend less time on satisfying the job requirement. This also highlights that activities, in general, are poor measures of skills.
3.5  Applying the adoption theorems under CES–CET
@@ 4  Discussion and Relation to the Literature
Task-based labor-market models have a long history. In contrast to much of this literature, and in line with international trade theory (Dixit and Norman, 1980; Helpman and Krugman, 1985), we model what workers can do as a convex production possibility set. This allows workers to flexibly adjust the tasks they supply in response not only to job requirements but also to technical change. Here we discuss how our approach differs from earlier models and how these differences affect our understanding of AI adoption.
Our framework rests on two key assumptions: task bundling and worker flexibility.
Task bundling refers to models where jobs require multiple tasks that are not tradable outside the specific worker-firm relationship. Foundational skill-bundle models (Welch, 1969; Katz and Murphy, 1992; Lazear, 2009; Autor and Thompson, 2025) treat workers as endowed with multiple skills that firms combine with heterogeneous weights. Sorting models (Gathmann and Schoenberg, 2010; Lise and Postel-Vinay, 2020; Bárány et al., 2020; Bárány and Holzheu, 2026) study how workers with multidimensional skills match to jobs with multidimensional requirements. Acemoglu and Autor (2011); Fortin and Lemieux (2016); Cavounidis and Lang (2020); Benzell and Myers (2026) use such task bundling models to study recent waves of automation.
Worker flexibility refers to whether workers can reallocate effort across tasks within a job. Models with flexibility include the routine-biased literature (Autor et al., 2003; Acemoglu and Autor, 2011; Cortes, 2016; Böhm, 2020), the automation-threshold models (Acemoglu and Restrepo, 2018, 2020; Chen and Liang, 2026), expertise compression (Autor and Thompson, 2025), and most recent AI papers (Freund and Mann, 2025; Althoff and Reichardt, 2026; Gans and Goldfarb, 2026; Cavounidis et al., 2024; Cavounidis and Lang, 2020). Most models assume perfectly elastic reallocation across tasks, which, in our model, corresponds to $\gamma=\infty$. This assumption, when tasks are not bundled results in market task prices, which are neither worker- nor job-specific. If tasks are bundled, each worker’s shadow prices are pinned down by their relative skills rather than by match-specific marginal valuations, and thus are not job-specific.
Very few models work with finite worker flexibility among a set of bundled tasks. Among the papers we survey, only Cavounidis and Lang (2020) allow workers to reallocate across tasks at a positive but finite cost, leading to the same CET production possibilities set as in our model. Their focus, however, is dynamic skill formation rather than technology adoption.
Our paper gives a general characterization of the technology adoption problem for any degree of worker and job flexibility, thereby naturally embedding endowment models ($\gamma=0$), market-based task models ($\sigma=\infty$) and any intermediate case. To the best of our knowledge, we are also the first to model AI adoption as a team production problem that makes the combined feasible set convex.
@@ 5  Conclusion
We model AI adoption as collaboration under a shared time and attention budget. A worker’s feasible task combinations form a convex production possibility set; an AI tool adds another such set with its own directional capabilities. Because using the tool requires supervision, the relevant object is the weighted sum of the worker and tool sets, with weights determined by time allocation. Adoption is therefore neither pure substitution nor pure augmentation: it is a team-production choice with an endogenous intensity.
This framing sharpens what can (and cannot) be predicted without context. The right sufficient statistic is the worker’s autarky (shadow) prices at their pre-AI optimum: they summarize what is scarce at the margin in the worker–job match and translate the tool’s multidimensional direction into a scalar value. Adoption occurs when the tool’s direction is valuable at these prices. Because shadow prices are about marginal valuations, they need not coincide with observed task allocations, skills, or job requirements taken separately.
Two implications follow for empirical work on GenAI in knowledge jobs. First, match-specificity is central: neither “which workers adopt” nor “which jobs are exposed” is well-defined in isolation because the same tool can be valuable in one match and irrelevant in another. Second, adoption has structured margins. As capabilities improve, adoption expands along an extensive margin (the set of adopted directions) and an intensive margin (partial versus all-in use), so small capability improvements near the threshold can generate large changes in usage while leaving other matches unchanged. These features help explain why scalar exposure indices can misclassify both adoption and gains when the technological frontier is jagged.
@@ A.5  CES-CET closed forms
$\frac{x_{i}}{x_{j}}=\left(\frac{\theta_{i}}{\theta_{j}}\right)^{\gamma/(\gamma+\sigma)}\left(\frac{s_{i}}{s_{j}}\right)^{\sigma/(\gamma+\sigma)}.$
@@
Time shares. From the allocation formula, $\omega_{i}^{A}=x_{i}^{A}/\sum_{j}x_{j}^{A}\propto s_{i}^{\sigma/(\gamma+\sigma)}\theta_{i}^{\gamma/(\gamma+\sigma)}$. Logarithmic differentiation gives $\partial\omega_{i}^{A}/\partial s_{i}=[\sigma/((\gamma+\sigma)s_{i})]\omega_{i}^{A}(1-\omega_{i}^{A})>0$.